I can not figure out the last steps on how the answer is attain
Please help show each step also how to enter it in the calculator for graphing solution.
3^(2x+2) =7^X
ln (3^2x+2) ==ln 7^x
(2x+2) ln (3) = x ln(7)
2x ln (3)= x ln 7-2 ln 3
2x ln (30-x(7)==-2 ln(3)
x(2 ln 3-ln 7) = -2 ln (3)
x=
Now I am stuck how do I finish it???? How is it enter into the calculator. Must know for my math class.
Any help, very appreciated.
thanks,
Jody

Answer by mukhopadhyay(490)

3^(2x+2) = 7^X
ln (3^2x+2) =ln 7^x
(2x+2) ln (3) = x ln(7)
2x ln (3)-x ln 7 = -2 ln 3
x[2 ln (3) - ln (7)] = -2 ln(3)
x = [2 ln (3) - ln (7) ]/ [-2 ln(3)]
x= ln (7)/[2 ln (3)] - 1